A 100W,250V lamp is connected in series with a 100W,200V lamp across a 250V supply mains. Calculate the current drawn from the supply mains and voltage across each lamp. Assume lamp resistance.
1
Expert's answer
2020-12-10T11:01:58-0500

Using nominal power and voltage of lamps find lamps' resintance:

"P=I\times U=\frac{U^2}{R}"

"R=\frac{U^2}{P}"

"R_{1}=\frac{250^2}{100}=625Ohm"

"R_{2}=\frac{200^2}{100}=400Ohm"

Current in circuit:

"I=\frac{U}{R_{1}+R_{2}}=\frac{250}{625+400}=0.244A"

Lamps' voltage:

"U=I\times R"

"U_{1}=I\times R_{1}=0.244\times 625=152.5V"

"U_{2}=I\times R_{2}=0.244\times 400=97.6V"